OpenStudy (anonymous):

Could someone please show how to find a Taylor polynomial for ln(x) centered around x=2? (Along with a brief methodology/explanation, if possible...)

OpenStudy (anonymous):

\[f(x)=\ln(x)\\ f'(x)=\frac{1}{x}=x^{-1}\\ f''(x)=-\frac{1}{x^2}=x^{-2}\\ f'''(x)=2x^{-3}\] first probably good to look for a pattern

OpenStudy (anonymous):

then we have to evaluate each of these at \(x=2\) and finally divide each by \(n!\) where \[c_n=\frac{f^{(n)}(2)}{n!}\]

OpenStudy (anonymous):

\[f(2)=\ln(2)\] for \(c_0\) no pattern there it is just the first term

OpenStudy (anonymous):

\[f'(2)=\frac{1}{2}\\ f''(2)=-\left(\frac{1}{2}\right)^2\] \[f'''(2)=2\times \left(\frac{1}{2}\right)^3\]

OpenStudy (anonymous):

pretty clear that it alternates also maybe it is clear that \[f^{(n)}{2}=(n-1)!\left(\frac{1}{2}\right)^n\]

OpenStudy (anonymous):

when you divide by \(n!\) you will be left with \[c_n=\frac{(-1)^n\left(\frac{1}{2}\right)^n}{n}\]

OpenStudy (anonymous):

guess i lost @james238867

OpenStudy (anonymous):

I'm still here. I'm actually understanding your explanation. Thanks!

OpenStudy (anonymous):

oh ok then i guess we are more or less done really it amounts to looking for a pattern, the less arithmetic you do the better, helps to see what the pattern is

OpenStudy (anonymous):

final answer will look something like \[\sum\frac{(-1)^n\left(\frac{1}{2}\right)^n(x-2)^n}{n}\]

OpenStudy (anonymous):

there are other ways to write it, you could put the \(2^n\) in the denominator

OpenStudy (anonymous):

also you should check that it is \((-1)^n\) or \((-1)^{n+1}\) i.e. check that the terms you want to be negative are in fact negative

OpenStudy (anonymous):

notice that you really don't want to do arithmetic as i said, and even though it is generally stupid to use the power rule for derivatives, you want to in this case \[f'(x)=x^{-1}\\ f''(x)=-x^{-2}\\ f'''(x)-2x^{-3}\\ f^{(4)}(x)=-3\times -2x^{-4}\] etc helps to get the pattern down

OpenStudy (anonymous):

I see. You've answered the question really completely. Thanks!

OpenStudy (anonymous):

yw